OpenStudy (anonymous):

Simplify (8 + 7i) + (2 − i). 10 + 6i 16 4 10 + 8i

OpenStudy (anonymous):

Simplify \[(8 +7i) + (2 - 1)\]

OpenStudy (kirbykirby):

\((a+bi)+(c+di)=(a+c)+i(b+d)\)

OpenStudy (kirbykirby):

wait is that -1 or -i ?

OpenStudy (anonymous):

oops, it - i

OpenStudy (anonymous):

wrote it wrong sorry

OpenStudy (anonymous):

\[(8 + 7i) + (2 - i) \]

OpenStudy (kirbykirby):

you can just think of this as like algebra, treating \(i\) as a variable like \(x\), just add like-terms :)

OpenStudy (anonymous):

(8 + 7i) + (2 - i) +i +i (8 + 8i) + 2

OpenStudy (anonymous):

then -2?

OpenStudy (kirbykirby):

o_o I'm not sure what you did there. It is much simpler than what you may think. Imagine the problem is written as \((8+7x)+(2-x)\) and imagine \(x=i\)

OpenStudy (anonymous):

so it's 10 + 6i

OpenStudy (kirbykirby):

\(\checkmark\_\checkmark\)

OpenStudy (anonymous):

i suck at math, that's what i did :D

OpenStudy (kirbykirby):

ya 10+6i is perfect ;) hehe.

OpenStudy (anonymous):

thanks!

OpenStudy (kirbykirby):

np!